OpenStudy (anonymous):

A spinner has numbers 1 to 4, a second spinner has numbers 1 to 3. The scores are added. What is the probability the total is a) 4 b) less than 5

OpenStudy (ribhu):

a) 1/4 = 0.25

OpenStudy (ribhu):

b)1/2

OpenStudy (aayushi.somani):

u can understand it better by making the pairs who add up to be 4 so here we get(1,3) , (2,2) , ( 3,1) so 3/12 (total no. of pairs)

OpenStudy (ribhu):

(1,1); (1,2) ; (1,3); (2,1); (2,2); (3,1) and total number of pairs would be 12 so the probability for b part is 6/12 = 1/2

OpenStudy (aayushi.somani):

yeah he's right for the second question

OpenStudy (anonymous):

I do not understand @aayushi.somani and @ribhu

OpenStudy (anonymous):

Please can you give a better explanation

OpenStudy (aayushi.somani):

ahh see the first spinner has 4 no.'s 1,2,3,4 and the second has 3 i.e 1,2,3 so we need to add the scores. in first part the sum should be exactly 4 (adding both the player's scores) so i created the pairs now in second part the sum should be less than 5 so i prepared the pairs accordingly.

OpenStudy (ribhu):

the numbers in the brackets indicate the number obtained on the respective spinner

OpenStudy (anonymous):

how come part b anser is 1/2

OpenStudy (anonymous):

@ribhu

OpenStudy (anonymous):

I understand that part a answer is 1/4

OpenStudy (aayushi.somani):

see make the pairs that add up to get a sum that is less than 5 it can be any no. among 1,2,3,4

OpenStudy (ribhu):

the answer to b part is as you have got 6 favourable outcomes and 12 total outcomes

OpenStudy (ribhu):

and the probability is favourable outcomes/ total outcomes

OpenStudy (anonymous):

but it only has number 1 -3 not 1-4

OpenStudy (ribhu):

since less than 5 was required and 1-4 would make the total as 5

OpenStudy (anonymous):

ok